OpenStudy (anonymous):

what is 2 =39/23 & 387

OpenStudy (anonymous):

What? That is not a legible equation

OpenStudy (anonymous):

I got it on my homework

OpenStudy (anonymous):

Well they must not have posted it like that

OpenStudy (anonymous):

Do you mean 2 = 39/23 x 387 Either wa, 2 =39/23 is false

OpenStudy (anonymous):

*way